A sports car accelerates in third gear from 48.3 km/h (about 30 mi/h) to 80.5km/h (about 50 mi/h) in 3.70s


A. What is the average acceleration of this car in mi/s?


B. If the car maintained this accelaration, how fast would it be moving one second later?


1
Expert's answer
2021-10-25T09:56:17-0400

1)

"a=\frac{v_2-v_1}t,"

"a=\frac{50-30}{3.7\cdot 3600}=1.5\cdot10^{-4}~\frac{mi}{s^2},"

2)

"v_3=v_2+at_1,"

"v_3=50+1.5\cdot 10^{-4}\cdot 3600\cdot 1=55.4~\frac{mi}h."